ANLA Quality List

HeOtto96 [1a] L. Hemmingsson and K. Otto, Analysis of semi-Toeplitz preconditioners for first-order PDEs, SIAM J. Sci. Comput., 17 (1996), pp. 47– 64.

Otto96 [1b] K. Otto, Analysis of preconditioners for hyperbolic partial differential equations, SIAM J. Numer. Anal., 33 (1996), pp. 2131–2165.

Hemm98 [1c] L. Hemmingsson, A semi-circulant preconditioner for the convectiondiffusion equation, Numer. Math., 81 (1998), pp. 211–248.

BrHoSu07 [2a] H. Brandén, S. Holmgren, and P. Sundqvist, Discrete fundamental solution preconditioning for hyperbolic systems of PDE, J. Sci. Comput., 30 (2007), published electronically in December 2005.

AxNey06 [2b] O. Axelsson and M. Neytcheva, Eigenvalue estimates for preconditioned saddle point matrices, Numer. Linear Algebra Appl., 13 (2006), pp. 339–360.

Larsson99 [3a] E. Larsson, A domain decomposition method for the Helmholtz equation in a multilayer domain, SIAM J. Sci. Comput., 20 (1999), pp. 1713–1731.

OttoLa99 [3b] K. Otto and E. Larsson, Iterative solution of the Helmholtz equation by a second-order method, SIAM J. Matrix Anal. Appl., 21 (1999), pp. 209–229.

BaNey05b [3c] E. Bängtsson and M. Neytcheva, Numerical simulations of glacial rebound using preconditioned iterative solution methods, Appl. Math., 50 (2005), pp. 183–201.